OpenStudy (anonymous):

solve the equation then check your solution 4x+7=-9

OpenStudy (anonymous):

i got4

OpenStudy (hba):

its -4

OpenStudy (hba):

4(-4)+7=-9 -16+7=-9 -9=-9